# ATI-CTLO: Adaptive Temporal Interval-based Continuous-Time LiDAR-Only Odometry

### v2

@@ III-C Non-Uniform Linear Interpolation Continuous-time LiDAR-Only Odometry baesed on Sliding Window
where the $\mathcal{F}_{ptpl}(\mathbf{X})$ is the point-to-plane registration errors, $\mathcal{F}_{vel}(\mathbf{X})$ is the constant velocity constraint. $\mathcal{F}_{m}(\mathbf{X})$ is the marginalization energy which keeps information of the removed variables when the sliding window slides. The error functions $\mathcal{F}_{ptpl}(\mathbf{X})$ and $\mathcal{F}_{vel}(\mathbf{X})$ of the whole sliding window consists of $M$ sub-error functions:
$\mathcal{F}_{ptpl}(\mathbf{X})$ $=\sum_{k=0}^{M-1}\mathcal{F}_{ptpl}\left(\mathbf{T}_{i+k},\mathbf{T}_{i+k+1}\right)$ (5a)
$\mathcal{F}_{vel}(\mathbf{X})$ $=\sum_{k=0}^{M-1}\mathcal{F}_{vel}\left(\mathbf{T}_{i+k},\mathbf{T}_{i+k+1}\right)$ (5b)
Every single sub-error function is associated with the linear segment $[\mathbf{T}_{i+k},\mathbf{T}_{i+k+1}]$ and the cloud segment $\mathbf{S}_{i+k}$. For the point-to-plane constraint:
$\mathcal{F}_{ptpl}(\mathbf{T}_{i+k},\mathbf{T}_{i+k+1})=\mathbf{e}_{p}(i+k)\mathbf{Q}^{-1}_{p}\mathbf{e}_{p}(i+k)$ (6a)
@@
where ${}^{L}\mathbf{p}_{j}$ is the $j$-th point of $N_{s}$ time-ordered points in $\mathbf{S}_{i+k}$, $\mathbf{T}_{j}$ transforms the point ${}^{L}\mathbf{p}_{j}$ to ${}^{W}\mathbf{p}_{j}$. ${}^{W}\mathbf{q}_{j}$ is the closest point to ${}^{W}\mathbf{p}_{j}$ in local map, and $\mathbf{n}_{q_{j}}$ denotes the normal vector of plane that includes ${}^{W}\mathbf{q}_{j}$. Then we calculate the error $\mathbf{e}_{p}(i+k)\sim\mathcal{N}(\mathbf{0},\mathbf{Q}_{p})$. Due to the constant velocity prior, $\mathbf{T}_{j}$ is calculated via linear interpolation as follows:
$\mathbf{T}_{j}$ $=\mathbf{T}_{i+k}\cdot\mathrm{Exp}(\boldsymbol{\alpha}_{j}\cdot\mathrm{Log}(\mathbf{T}_{i+k}^{-1}\mathbf{T}_{i+k+1}))$ (7a)
$\boldsymbol{\alpha}_{j}$ $=(t_{j}-t_{i+k})/(t_{i+k+1}-t_{i+k})$ (7b)
As for $\mathcal{F}_{vel}(\mathbf{T}_{i+k},\mathbf{T}_{i+k+1})$, we first calculate the kinematic error $\mathbf{e}_{v}(i+k)\sim\mathcal{N}(\mathbf{0},\mathbf{Q}_{v})$. Due to the varying intervals between adjacent linear segments, a scaling factor $\boldsymbol{\gamma}_{i+k}$ must be incorporated into the constraint calculations:
$\mathcal{F}_{vel}(\mathbf{T}_{i+k},\mathbf{T}_{i+k+1})=\mathbf{e}^{\top}_{v}(i+k)\mathbf{Q}^{-1}_{v}\mathbf{e}_{v}(i+k)$ (8a)
@@
When the sliding window slides forward, old control nodes $\mathbf{X}_{m}$ are removed through marginalization, While the corresponding cloud segments $\mathbf{S}_{m}$ are added to local map. The energy $\mathcal{F}_{m}$ keeps information from $\mathbf{X}_{m}$, which is commonly stored as:
$\mathcal{F}_{m}=\frac{1}{2}(\mathbf{X}_{m}-\mathbf{X}_{m}^{0})^{\top}$ $H_{m}(\mathbf{X}_{m}-\mathbf{X}_{m}^{0})+b^{\top}_{m}(\mathbf{X}_{m}-\mathbf{X}_{m}^{0})$ (9a)
@@
where the $H_{m}$,$b_{m}$ is marginalization priors, the $\mathbf{X}_{m}^{0}$ is the linearization point of the removed nodes, which is applied for calculating the Schur complement $\tilde{H}$ and corresponding vector $\tilde{b}$. From the $H_{m}$, $b_{m}$, we can derive $J_{m}$ and $r_{m}$ as (9b), please refer to [14] for more details.
Figure 4: Degeneracy Detection: Each column of the information matrix $\mathbf{I}$ represents the projection of environmental information in the corresponding pose direction. In the contribution evaluation, we calculate the contribution of each column $\mathbf{I}_{i}$ and use $l_{c}$ and $l_{s}$ to count the number of information elements $\mathbf{I}_{(i,j)}$ with varying levels of contribution. Then different thresholds $\kappa_{1},\kappa_{2},\kappa_{3}$ are used to classify the contributions into various levels, eventually the localizability of each direction is divided into three levels: {localizable, partially localizable, nonlocalizable}. Degeneracy Management: we enhance the number of effective information elements by increasing the sampling rate (decreasing $D_{s}$) for partially localizable direction. In nonlocalizable case, subsequent cloud segment $\mathbf{S}_{i+1}$ is merged to gather more valid information.
@@ III-D1 Degeneracy Detection
As the linear segments $\mathbf{S}_{i}$ are obtained by slicing the LiDAR scans $\mathbf{L}$, they accumulate fewer LiDAR points compared to discrete-time methods, which increases the likelihood of degeneracy during optimization. In this paper, a classification of degeneracy levels was conducted using geometric feature-based methods, with specific handling approaches applied to each level. This approach evaluates how much a information pair $(^{L}\mathbf{p},^{L}\mathbf{n})$ contributes the localizability towards principle direcitons in optimization. Similar to [31], the normal vector ${}^{L}\mathbf{n}$ and torque $\boldsymbol{\tau}=^{L}\mathbf{p}\times^{L}\mathbf{n}$ in LiDAR frame respectively provide localizability contributions toward translational $\mathbf{v}_{t}$ and rotational $\mathbf{v}_{r}$ directions, which are the eigenvectors of Hessian matrix of optimization. When the normal vector ${}^{L}\mathbf{n}$ or torque $\boldsymbol{\tau}$ aligns with the principal directions of optimization (${}^{L}\mathbf{n}\cdot\mathbf{v}_{t}>0$ or $\boldsymbol{\tau}\cdot\mathbf{v}_{r}>0$), it is deems to provide a localizability contribution. Due to the point ${}^{L}\mathbf{p}$ can be at long distance, the $\boldsymbol{\tau}$ need to be normalized first, so the information matrix is as:
@@
In this letter, the sliding window is used to optimize several linear segments at once, but it is unnecessary to evaluate all information pairs within the window as it can lead to repeated assessments of the same linear segment. As Fig. 2 shows, for each optimization, we only focus on the latest cloud segment $\mathbf{S}_{i+2}$. We perform SVD decomposition of the Hessian Matrix of $\mathbf{T}_{i+3}$ to obtain the principal direction matrices of rotation $\mathbf{V}_{r}$ and translation $\mathbf{V}_{t}$. By projecting the information matrices $\mathbf{F}$ onto the main optimization directions $\mathbf{V}$, we can determine the contributions in each respective direction:
@@ III-D2 Degeneracy Management
In discrete-time methods, the nonloclizable directions are excluded from the optimization, and prior motions values are used instead. For partially localizable directions, the constraints are calculated by resampling the available information instead.
In this letter, degeneracy may arise from the segmentation of the point cloud, which reduces the number of constraint points. Thus, we must first exclude degeneracy caused by this factor. As Fig. 4 shows, if a particular direction is nonlocalizable, we will merge the $\mathbf{S}_{i}$ with the next segment $\mathbf{S}_{i+1}$ and then perform degeneracy detection again. To prevent unlimited merging, we stop the process when the temporal interval $\Delta t_{S_{i}}$ of segment $\mathbf{S}_{i}$ exceeds $0.1$ seconds. If a direction is partially localizable, it indicates that $\mathbf{S}_{i}$ contains useful information. We then halve the voxel size for downsampling $\mathbf{S}_{i}$ to obtain more available information pairs. Similarly, when the voxel size falls below 0.1, we move directly to the optimization.
TABLE I: ATE of LiDAR-Only Odometry on M2DGR Dataset
Approach street_03 street_07 room_01 room_dark_06 gate_01 hall_02 door_02 circle_01 average
LeGO_LOAM 0.597 10.584 0.136 0.305 0.328 0.239 0.187 1.529 1.738
KISS-ICP 0.131 0.400 0.219 0.339 0.141 0.276 0.263 1.499 0.408
CT-ICP 0.174 x 0.144 0.293 0.159 0.259 0.211 0.357 0.228
Traj-LO 0.173 0.249 0.148 0.293 0.151 0.257 0.196 0.357 0.228
ours 0.170 0.212 0.132 0.284 0.151 0.251 0.185 0.354 0.186
TABLE II: NTU VIRAL Dataset
@@
FLOAM L1 4.486 8.328 1.133 1.447 1.292 1.498 0.976 0.2010 1.079 10.775 4.637 2.218 2.354 2.249 1.566 x x x
KISS-ICP L1 2.220 1.570 1.014 0.628 1.500 1.272 0.917 1.312 1.030 3.663 1.970 2.382 2.305 2.405 0.799 8.493 x 5.454
CT-ICP L1 7.763 0.125 11.171 0.100 0.101 0.073 x 0.084 1.545 x 0.081 0.086 0.073 0.071 0.045 x x x
Traj-LO L1 0.055 0.039 0.035 0.047 0.052 0.050 0.048 0.039 0.039 0.050 0.058 0.057 0.505 0.607 0.101 0.121 x 0.103
@@ IV EXPERIMENTS
We validated the effectiveness of the ATI-CTLO algorithm on datasets from various platforms, including UGV dataset M2DGR [32] , UAV dataset NTU VIRAL [33] and our own quadruped robot dataset. In comparisons with the state-of-the-art LiDAR odometry methods, including discrete-time methods LeGO-LOAM and KISS-ICP, and fixed-interval continuous-time odometry algorithms like CT-ICP and Traj-LO, our method ATI-CTLO demonstrated an ability to handle more challenging environments while improving accuracy by around 10% in ground robots.
In this letter, we maintained the same parameters setting for all experiments. In the PCA pre-assessment module, when the change in eigenvalue is below $\mathcal{K}_{val}=0.1$ and the angle $\boldsymbol{\varphi}_{max}$ exceeds $\mathcal{K}_{vec}=10$, we halve the temporal interval ($\alpha=0.5$). In the degeneracy management module, we use $\kappa_{1}=125,\kappa_{2}=50,\kappa_{3}=15$ to classify the levels of degeneracy. All experiments were conducted on a Intel i7-12700H CPU.
IV-A Dataset Result
@@ IV-A1 M2DGR Dataset
Figure 5: Results of LiDAR odometry on street_07 sequence of M2DGR. (a) Our method, the yellow box marks the position where CT-ICP drifts. (b) CT-ICP drifts. (c) Trajectory results aligned with the ground truth. (d) A zoomed-in view, where orange boxes indicates increased temporal intervals between control nodes due to degeneracy management, and black boxes indicates reduced intervals due to PCA-assessed motion state changes
We first validate our algorithm’s effectiveness under vehicle conditions. The M2DGR dataset, collected by a UGV (Unmanned Ground Vehicle), provided data from various environments. We conduct tests in each environment and compute the RMSE of ATE error for the trajectories. The results are shown in Table I, showing that our method achieved an approximate 10% improvement in average accuracy compared to the state-of-the-art (SOTA) LiDAR-Only methods (excluding the street_07 sequence for average accuracy of CT-ICP). In sequences such as street_03, gate_01, and hall_02, where the vehicle’s movement is relatively smooth, discrete-time methods outperform continuous-time methods in terms of accuracy. However, in sequences with aggressive turns, such as street_07 and circle_01, the errors of discrete-time methods increase significantly. Across all sequences, continuous-time methods maintain more stable accuracy, and our method consistently achieve the highest accuracy among continuous-time methods. This improvement is due to our dynamic temporal interval adjustment strategy.
As seen in Fig. 5(d), our method continuously adjusts the time intervals, resulting in a more accurate alignment with the ground truth. In contrast, CT-ICP experiences significant drift during execution (Fig. 5(b)), leading to automatic termination.

### v3

@@ III-C Non-Uniform Linear Interpolation Continuous-time LiDAR-Only Odometry Based on Sliding Window
where the $\mathcal{F}_{ptpl}(\mathbf{X})$ is the point-to-plane constraint, $\mathcal{F}_{vel}(\mathbf{X})$ is the constant velocity constraint. $\mathcal{F}_{m}(\mathbf{X})$ is the marginalization energy that keeps information on the removed variables when the sliding window moves. The error functions $\mathcal{F}_{ptpl}(\mathbf{X})$ and $\mathcal{F}_{vel}(\mathbf{X})$ of the whole sliding window consists of $M$ sub-error functions, which correspond to $M$ cloud segments $\{\mathbf{S}_{i},\mathbf{S}_{i+1},\cdots,\mathbf{S}_{i+M-1}\}$:
$\mathcal{F}_{ptpl}(\mathbf{X})$ $=\sum_{k=0}^{M-1}{\mathcal{F}_{sub\_ptpl}}\left(\mathbf{T}_{i+k},\mathbf{T}_{i+k+1}\right)$ (5a)
$\mathcal{F}_{vel}(\mathbf{X})$ $=\sum_{k=0}^{M-1}{\mathcal{F}_{sub\_vel}}\left(\mathbf{T}_{i+k},\mathbf{T}_{i+k+1}\right)$ (5b)
Every single sub-error function is associated with the linear segment $[\mathbf{T}_{i+k},\mathbf{T}_{i+k+1}]$ and the cloud segment $\mathbf{S}_{i+k}$. For the point-to-plane constraint, as described in [8] :
${\mathcal{F}_{sub\_ptpl}}(\mathbf{T}_{i+k},\mathbf{T}_{i+k+1})=\mathbf{e}_{p}(i+k)\mathbf{Q}^{-1}_{p}\mathbf{e}_{p}(i+k)$ (6a)
@@
where the $\mathcal{F}_{sub\_ptpl}$ is derived from the point-to-plane residual $e_{p}$. Due to the constant velocity prior, $\mathbf{T}_{j}$ is calculated via linear interpolation.
As for $\mathcal{F}_{sub\_vel}(\mathbf{T}_{i+k},\mathbf{T}_{i+k+1})$, differing from fixed time interval methods, a scaling factor $\boldsymbol{\gamma}_{i+k}$ is incorporated into the velocity constraint calculations:
${\mathcal{F}_{sub\_vel}}(\mathbf{T}_{i+k},\mathbf{T}_{i+k+1})=\mathbf{e}^{\top}_{v}(i+k)\mathbf{Q}^{-1}_{v}\mathbf{e}_{v}(i+k)$ (7a)
@@
As the sliding window moves backward, the cloud segment $\mathbf{S}_{i}$ is removed and control node $\mathbf{T}_{i}$ is marginalized, while $\mathbf{T}_{i+1}$, which is still relevant to $\mathbf{S}_{i+1}$, is retained. Hence, $\mathbf{X}_{m}$ also includes $\mathbf{T}_{i+1}$ coming from $\mathbf{S}_{i}$, with this relationship represented by $\mathbf{X}_{m}=marg(\mathbf{X})$. The energy $\mathcal{F}_{m}(\mathbf{X})$ keeps information from $\mathbf{X}_{m}$, which is commonly stored as:
$\mathcal{F}_{m}(\mathbf{X})=\frac{1}{2}(\mathbf{X}_{m}-\mathbf{X}_{m}^{0})^{\top}$ $H_{m}(\mathbf{X}_{m}-\mathbf{X}_{m}^{0})+b^{\top}_{m}(\mathbf{X}_{m}-\mathbf{X}_{m}^{0})$ (8a)
@@
where the $H_{m}$,$b_{m}$ is marginalization priors, the $\mathbf{X}_{m}^{0}$ is the linearization point, which is applied for calculating the Schur complement $\tilde{H}$ and corresponding vector $\tilde{b}$. From the $H_{m}$, $b_{m}$, we can derive $J_{m}$ and $r_{m}$ as Equ. (8b), please refer to [14] for more details.
Figure 4: Degeneracy Detection: Each column $\mathbf{I}_{j}$ of the contribution matrix $\mathbf{I}$ represents the projection of environmental information in the corresponding principal direction. In the contribution evaluation, we calculate the contribution of each column $\mathbf{I}_{j}$ and use $l_{c}$ and $l_{s}$ to count the number of information elements $\mathbf{I}_{(i,j)}$ with varying levels of contribution. Then different thresholds $\kappa_{1},\kappa_{2},\kappa_{3}$ are used to classify the contributions into various levels, eventually the localizability of each direction is divided into three levels: {localizable, partially localizable, nonlocalizable}. Degeneracy Management: We enhance the number of effective information elements by increasing the sampling rate (decreasing $D_{s}$) for partially localizable cases. In nonlocalizable cases, subsequent cloud segment $\mathbf{S}_{i+3}$ is merged to gather more valid information.
@@ III-D1 Degeneracy Detection
As the linear segments $\mathbf{S}_{i}$ are obtained by slicing the LiDAR scans $\mathbf{L}$, they accumulate fewer LiDAR points compared to discrete-time methods, which increases the likelihood of degeneracy during optimization. In this letter, a classification of degeneracy levels is conducted using geometric feature-based methods, with specific handling approaches applied to each level. As described in [31][32], this approach evaluates the contribution of an information pair $(^{L}\mathbf{p},^{L}\mathbf{n})$ to the localizability in the principal directions of optimization by calculating the normal vector ${}^{L}\mathbf{n}$ and torque $\boldsymbol{\tau}=^{L}\mathbf{p}\times^{L}\mathbf{n}$. So the information matrix is defined as:
@@
Differing from [31], the sliding window is used to optimize several multiple control nodes at once, with each cloud segment $\mathbf{S}_{i}$ corresponding to two control nodes $T_{i}$ and $T_{i+1}$. Performing degeneracy detection on all variables within the sliding window would significantly reduce the algorithms real-time performance. In practice, only the latest point cloud segment within the sliding window is considered during implementation. For example, in Fig 2, only the information from $\mathbf{S}_{i+2}$ is projected onto the optimization principal directions $\mathbf{V}_{r}$ and $\mathbf{V}_{t}$ corresponding to $\mathbf{T}_{i+3}$ to obtain the contribution matrix:
@@ III-D2 Degeneracy Management
In discrete-time methods, the nonloclizable directions are excluded from the optimization, and prior motion values are used instead. For partially localizable directions, the constraints are calculated by resampling the available information instead.
In this letter, degeneracy may arise from the segmentation of the point cloud, which reduces the number of constraint points. Thus, we must first exclude degeneracy caused by this factor. As Fig. 2 shows, if a particular direction is nonlocalizable, we will merge the $\mathbf{S}_{i+2}$ with the next segment $\mathbf{S}_{i+3}$ and then perform degeneracy detection again. To prevent unlimited merging, we stop the process when the temporal interval $\Delta t$ of segment $\mathbf{S}_{i+2}$ exceeds $0.1s$. If a direction is partially localizable, it indicates that $\mathbf{S}_{i+2}$ contains useful information. We then halve the voxel size for downsampling $\mathbf{S}_{i+2}$ to obtain more available information pairs. Similarly, when the voxel size falls below 0.1, we move directly to the optimization.
TABLE I: RMSE of ATE (m) / RPE (m) on M2DGR Dataset
Approach room_01 room_dark_06 gate_01 hall_02 door_02 circle_01 street_03 street_07 Average
(27.1m) (72.5m) (139.2m) (109.4m) (57.3m) (316.5m) (423.9m) (1104.6m)
LeGO-LOAM 0.136 / 0.203 0.305 / 0.217 0.328 / 0.482 0.239 / 2.693 0.187 / 2.364 1.529 / 0.710 0.597 / 0.802 10.584 / 0.169 1.738 / 0.955
KISS-ICP 0.219 / 0.114 0.339 / 0.125 0.141 / 0.147 0.276 / 0.809 0.263 / 0.406 1.499 / 0.254 0.131 / 0.256 0.400 / 0.164 0.408 / 0.284
CT-ICP 0.144 / 0.089 0.293 / 0.091 0.159 / 0.152 0.259 / 0.808 0.211 / 0.404 0.357 / 0.252 0.174 / 0.259 ✗ / ✗ 0.228 / 0.294
Traj-LO 0.148 / 0.043 0.293 / 0.036 0.151 / 0.056 0.257 / 0.362 0.196 / 0.170 0.357 / 0.103 0.173 / 0.126 0.249 / 0.065 0.228 / 0.120
Ours 0.132 / 0.040 0.284 / 0.037 0.151 / 0.059 0.251 / 0.364 0.185 / 0.157 0.354 / 0.099 0.170 / 0.101 0.212 / 0.064 0.217 / 0.115
  • 1
    ’✗’ denotes significant error, occurring when the estimated distance or orientation between consecutive control nodes exceeding $5m$ or $30^{o}$ [8].   ’(*m)’ denotes the length of the sequence.
TABLE II: RMSE of ATE (m) on NTU VIRAL Dataset
@@
(237.0m) (171.1m) (127.8m) (160.3m) (249.1m) (315.5m) (202.3m) (183.5m) (198.5m) (321.2m) (200.0m) (304.5m) (252.3m) (139.6m) (123.8m) (545.0m) (697.9m) (312.4m)
F-LOAM L1 4.486 8.328 1.133 1.447 1.292 1.498 0.976 0.2010 1.079 10.775 4.637 2.218 2.354 2.249 1.566
KISS-ICP L1 2.220 1.570 1.014 0.628 1.500 1.272 0.917 1.312 1.030 3.663 1.970 2.382 2.305 2.405 0.799 8.493 5.454
CT-ICP L1 7.763 0.125 11.171 0.100 0.101 0.073 0.084 1.545 0.081 0.086 0.073 0.071 0.045
Traj-LO L1 0.055 0.039 0.035 0.047 0.052 0.050 0.048 0.039 0.039 0.050 0.058 0.057 0.505 0.607 0.101 0.121 0.103
@@
  • 1
    ’✗’ denotes significant error, as defined in the same way as in Table I.   ’(*m)’ denotes the length of the sequence.
@@ IV EXPERIMENTS
We validated the effectiveness of the ATI-CTLO algorithm on datasets from various platforms, including the UGV dataset M2DGR [33] , the UAV dataset NTU VIRAL [34] , the handheld dataset Newer College Dataset  [35] and our own quadruped robot dataset. In comparisons with the SOTA LiDAR odometry methods, including discrete-time methods such as LeGO-LOAM [17] and KISS-ICP [3], and fixed-interval continuous-time odometry algorithms like CT-ICP [8] and Traj-LO [14], our method ATI-CTLO demonstrated an ability to handle more challenging environments.
In this letter, we maintained the same parameters setting for all experiments. In the PCA pre-assessment module, when the change in eigenvalue is below $\mathcal{K}_{val}=0.1$ and the angle $\boldsymbol{\varphi}_{max}$ exceeds $\mathcal{K}_{vec}=10$, we halve the temporal interval ($\alpha=0.5$). In the degeneracy management module, we used $\kappa_{1}=125,\kappa_{2}=50,\kappa_{3}=15$ to classify the levels of degeneracy. For the optimization process, the initial temporal interval $\Delta t$ was set to $0.04s$, and the capacity of sliding window $M$ was set to $3$. All experiments were conducted on an Intel i7-12700H CPU.
IV-A Dataset Results
@@ IV-A1 M2DGR Dataset
Figure 5: Results of LiDAR odometry on street_07 sequence of M2DGR. (a) Our method, the white box marks the position where CT-ICP drifts. (b) CT-ICP shows significant errors at the marked position in (a) . (c) Trajectory results aligned with the ground truth. (d) A zoomed-in view, where orange boxes indicate increased temporal intervals between control nodes due to degeneracy management, and black boxes indicate reduced intervals due to PCA-assessed motion state changes.
We first validate our algorithm’s effectiveness under vehicle conditions. The M2DGR dataset, collected by a unmanned ground vehicle (UGV). We conduct tests in various environments and compute the Root Mean Squared Error (RMSE) of Absolute Trajectory Error (ATE) and Relative Pose Error (RPE) for the trajectories. As shown in Table I, while the average improvement in accuracy on the UGV is modest (5%, calculated from ATE), our method achieves a remarkable improvement (14%, calculated from ATE) in the highly dynamic sequence (street_07). As seen in Fig. 5(d), our method continuously adjusts the temporal intervals, resulting in a more accurate alignment with the ground truth.